OpenStudy (mach_15):

What is the length of the major axis? a) 18 units b) 12 units c) 20 units d) 50 units

OpenStudy (mach_15):

\[\frac{ (x+2)^2 }{ 100 }+\frac{ (y+7)^2 }{ 36 }=1\]

OpenStudy (loser66):

100 and 36, which one is larger?

OpenStudy (mach_15):

100

OpenStudy (loser66):

Ok, then \(100=?^2\)

OpenStudy (mach_15):

10

OpenStudy (loser66):

yeah!! that's the answer

OpenStudy (loser66):

don't get?

OpenStudy (loser66):



OpenStudy (loser66):

Thanks for supporting @mathmate @mach_15 I am sorry for my wrong answer, it is 2*10 = 20

OpenStudy (loser66):



OpenStudy (mach_15):

Oh okay. thanks!